OpenStudy (anonymous):

How do I find the angle between the sphere x^2+y^2+z^2-9=0 and the paraboloid z-x^2-y^2+3=0 at point (2,-1,2) ?

OpenStudy (anonymous):

@sidsiddhartha Could you help with this please? (:

OpenStudy (amistre64):

tangent planes might be a start

OpenStudy (anonymous):

Should I derive @amistre64

OpenStudy (amistre64):

do you recall what a gradient is?

OpenStudy (anonymous):

f(x,y)+fx(x,y)+fy(x,y)?

OpenStudy (amistre64):

that sounds about right, the gradiant tells us the normal of the tangent plane at a given point ... it is those vectors that you need to be comparing

OpenStudy (amistre64):

find f wrt.x and f wrt.y and f wrt.z

OpenStudy (anonymous):

for both of them?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

uh so for the sphere <2x,2y,2z> and paraboloid <-2x,-2y,1>

OpenStudy (amistre64):

yep let: f = x^2+y^2+z^2-9 and g = z-x^2-y^2+3 find the partials for f and g, to define the normal vectors to their tangent planes and then we can determine the angles between them

OpenStudy (amistre64):

apply the xyz from the point into the gradients to define them now

OpenStudy (sidsiddhartha):

\[\nabla \phi_1=2x \hat{i}+2y \hat{j}+2z \hat{k}\\and\\ \nabla \phi_2=-2x \hat{i}-2y \hat{j}+\hat{k}\] like this

OpenStudy (anonymous):

uh right and then?

OpenStudy (sidsiddhartha):

\[now~\angle~\between~them\\cos \theta=\frac{ \nabla \phi_1. \nabla \phi_2 }{ |\nabla \phi_1|.|\nabla \phi_2|}\]

OpenStudy (amistre64):

recall the dotproduct trig definition |u| |v| cos(a) = u.v that will help us define the angle between the normals

OpenStudy (anonymous):

Gotcha thank you guys so much

OpenStudy (sidsiddhartha):

yes simple dot product